OpenStudy (anonymous):

Herman is making a small triangular table, in the shape of a right triangle. If two of the sides will measure about 25 inches each, how many inches should the longest side (c) measure? Use the appropriate formula and round your answer to the nearest tenth. A. 50 in B. 40.5 in C. 35.4 in Herman is making a small triangular table, in the shape of a right triangle. If two of the sides will measure about 25 inches each, how many inches should the longest side (c) measure? Use the appropriate formula and round your answer to the nearest tenth. A. 50 in B. 40.5 in C. 35.4 in @Mathematics

OpenStudy (anonymous):

alright so we have two sides with the same length and it's a right triangle, so that means the two sides we are given are legs. so we have (25)^2+(25)^2=L^2 where L is our length....or L= 25*sqrt 2 which is roughly 35.4inches :) so C